OpenStudy (anonymous):

use a commutative law to find an expression equivalent to 100+a

mathslover (mathslover):

a + 100

Parth (parthkohli):

Commutative law says a + b = b + a Similarly, \(\Large \color{MidnightBlue}{\Rightarrow 100 + a = a + 100 }\)